OpenStudy (anonymous):

you roll a die. What is the probability that you roll either a 5 or 6?

OpenStudy (anonymous):

2/6??? =1/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you sure???????? lol

OpenStudy (anonymous):

yes can you help me

OpenStudy (anonymous):

ok ok hold on

OpenStudy (anonymous):

p(5n6)=1/6*1/6=1/36

OpenStudy (anonymous):

why is it 1/36???

OpenStudy (anonymous):

sorry it is 1/3 as or mean plus

OpenStudy (anonymous):

WRONG!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

thank you